OpenStudy (anonymous):

In a recent golf match, Desmond’s score was 3 less than George’s score. Their combined scores totaled 187. Let d represent Desmond's score and g represent George's score. Which pair of equations can be used to determine their scores? d + g = 187 d – g = 3 d = g – 3 d + g = 187 g = d – 3 d + g = 187 d + g = 3 187 + d = g

Directrix (directrix):

What would be the equation for this: Their combined scores totaled 187 @CrazyCountryGirl

Directrix (directrix):

Use those d and g variables from the posted question, okay?

Directrix (directrix):

Yo! Help me out here, please.

Directrix (directrix):

What is the sum of d and g? That is what I am asking.

Directrix (directrix):

Yo is a word. Click here: http://www.merriam-webster.com/dictionary/yo

Directrix (directrix):

Their combined scores totaled 187 d + g = 187

Directrix (directrix):

Desmond’s score was 3 less than George’s score. d = ? -3

OpenStudy (anonymous):

@Directrix Okay, can we clear all the comments and start this over? I'm really confused right now.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

We have this: d + g = 187 That is correct We need one more: If you fill in the letter here, we will be finished. Desmond’s score was 3 less than George’s score. d is Desmond; g is George d = ? -3 --> all you have to do is to give the letter of the guy's name that goes in the question mark position. Then, we will be finished.

OpenStudy (anonymous):

g?

Directrix (directrix):

yes d = g -3 d + g = 187

Directrix (directrix):

I see those in the options.